OpenStudy (anonymous):

Solve for x: 1 over 2 plus 4 over 2 x equals quantity x plus 4 over 10.

OpenStudy (anonymous):

Will medal please help!

OpenStudy (anonymous):

can you make it an equation please

OpenStudy (anonymous):

1/2 + 4/2x= x+4/10

OpenStudy (anonymous):

\[\frac{ 1 }{ 2} + \frac{ 4 }{ 2 }\left( x \right)=\frac{ x+4 }{ 10 }\] ?

OpenStudy (anonymous):

no the x is with the 2 on the 2nd fraction

OpenStudy (anonymous):

Ok so what do you think you would do 1st?

OpenStudy (anonymous):

find the common denominator which is 10x

OpenStudy (anonymous):

Yeah so do that and tell me what you get

OpenStudy (anonymous):

5x\10x+ 20\10x= x2+4x\10x?

OpenStudy (anonymous):

@taylorz115820

OpenStudy (anonymous):

Yep I am pretty sure that is right Now what would you do?

OpenStudy (anonymous):

combine the left side

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

Tell me what you get

OpenStudy (anonymous):

The answer I got was 5

OpenStudy (anonymous):

as in x=5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so you can check by plugging 5 into the original equation and checking to see if they equal each other

OpenStudy (anonymous):

So if they equal each other then you have done the problem right

OpenStudy (anonymous):

Did it work?